OpenStudy (anonymous):

Which expression is equivalent to the area of metal sheet required to make this square shaped traffic sign?

OpenStudy (anonymous):



OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

no i need help. i dont know where to start

OpenStudy (skullpatrol):

Start with the formula for the area of a square.

OpenStudy (anonymous):

lxwxh?

OpenStudy (skullpatrol):

That is the formula for the volume of a rectangular box, all we need here is A=lxw

OpenStudy (anonymous):

so (2x-1)^2?

OpenStudy (anonymous):

or 2(2x-1)

OpenStudy (skullpatrol):

$$(2x-1)^2$$ is the correct answer because A = l*w and l=w=2x-1...good work :)

OpenStudy (anonymous):

so what do i do after that? would it turn into 4x^2=1? here are my choices..... 4x2 - 1 4x2 + 1 4x2 + 4x - 1 4x2 - 4x + 1

OpenStudy (skullpatrol):

$$A=l*w=(2x-1)(2x-1)=(2x-1)^2$$ Use the FOIL method.

OpenStudy (skullpatrol):

$$4x^2 - 1$$ $$4x^2 + 1$$ $$4x^2 + 4x - 1$$ $$4x^2 - 4x + 1$$

OpenStudy (skullpatrol):